OpenStudy (anonymous):

rationalize the denominator....... 1/(sqrt 5)(sqrt 2)

myininaya (myininaya):

\[\frac{1}{\sqrt{5}\sqrt{2}}\] this right?

OpenStudy (anonymous):

no.. it was a wrong question.... i forgot to put a + between sqrt 5 and sqrt 2

myininaya (myininaya):

\[\frac{1}{\sqrt{5}+\sqrt{2}}*\frac{\sqrt{5}-\sqrt{2}}{\sqrt{5}-\sqrt{2}}=\frac{\sqrt{5}-\sqrt{2}}{5-2}=\frac{\sqrt{5}-\sqrt{2}}{3}\]